# An Adaptive Placement and Parallelism Framework for Accelerating RLHF Training

### 19 Dec 2023 v1

@@ VI-B1 Comparing with DeepSpeed-Chat
Separation vs Flattening Strategy. The experiments for the Separation strategy are specifically conducted in large-scale training scenarios, such as training 33B and 65B models on $4\times 8$ and $16\times 8$ GPU devices. However, this strategy is not well-suited for small training scenarios with a limited number of GPUs and smaller models. The reason is that the Separation strategy involves adding a replication of the Actor and Critic models exclusively for the prediction stage, which results in increased memory redundancy.
However, the benefits of the Separation strategy are significant, as demonstrated in Table II. Our Separation strategy outperforms the Flattening strategy used in DeepSpeed-Chat by factors ranging from 4$\times$ to 11$\times$. For instance, in 33B model training using 4$\times$8 GPU devices, the throughput increases from 0.93 to 4.76 samples per second, resulting in a speedup ratio of 4$\times$. The speedup ratio further increases to 7$\times$ when the model size is increased to 65B. Additionally, the Separation strategy achieves even higher speedup when scaling out to $16\times 8$ GPU devices, reaching a speedup ratio of 11$\times$ compared to the baseline using DeepSpeed-Chat. The contribution comes from that the Separation strategy decouples the training and generation stages of Actor model, which is not possible for the Flattening strategy. This decoupling allows the usage of intra-node tensor parallelism for the generation stages as discussed in section V-A. For example, in our training of 65B model using $16\times 8$ GPUs, $3\times 8$ devices are exclusively allocated for the Pred Actor model. In this case, we apply tensor parallelism within the same node while replicating it across three machines. This approach significantly reduces inter-node communications compared to frequent inter-node AllGather and Scatter communications in ZeRO-3, which are required in the 33B and 65B model training for the Flattening strategy.
@@
Heterogeneous Separation vs DeepSpeed-Chat. We also evaluate the Heterogeneous Separation strategy in a heterogeneous GPU cluster. The cluster consists of 8 A100 GPUs and an additional 8 spare V100 GPUs with 32GB memory. These heterogeneous V100 GPUs can not be utilized under the Flattening strategy. We specifically conduct experiments using large models, such as the 33B model. Under our heterogeneous Separation strategy, the Reference and Reward models are placed on a separate node with 4 V100 GPUs respectively. We enable the intra-node tensor parallelism for both models and allocate two A100 GPUs to two Pred Actor replicas exclusively. This configuration aims to accelerate the generation speed and improve the overall training speed.
As shown in Table III, introducing the idle V100 devices under the heterogeneous Separation strategy results in a speedup of 14.49%. This acceleration increases the throughput from 0.69 to 0.79 samples per second. Although the acceleration may not be considerable, this strategy explores a novel solution that allows RLHF training to utilize idle heterogeneous resources to enhance the overall throughput. It re-balances the inference workloads among different GPU devices by simply adjusting the batch size, without requiring extensive consideration of software and hardware heterogeneity among different types of GPUs.
VI-B2 Comparing with trlX
In this subsection, we compare our APP framework against trlX, which employs a variant of the Flattening strategy, in the AC-Share scenario. We evaluate our implementation of Flattening and Interleaving strategies in the APP framework in terms of throughput under various computing resources. However, before presenting the experimental results, there are three important notes to mention:
i) To align the number of models with trlX for a fair comparison, our Flattening and Interleave strategies only consider the Actor, Reference, and Reward models, excluding the Critic model in terms of placement. The reason is that the Actor and Critic models share most of the parameters in AC-share structure. The modifications required are minimal, thanks to the modular design of our model placement abstraction.
@@ VI-B2 Comparing with trlX
TABLE IV: Comparing our Flattening strategy with trlX on $1$ machine with $8$ GPU devices. Two kinds of trlX implementation are employed.
@@
Flattening 10.48 2.9
ON trlX (stand-alone RM) 5.92 2.87
@@
Flattening 25.58 11.53
TABLE V: Comparing our Flattening and Interleaving strategy with trlX on varying numbers of machines, each with $8$ GPU devices.
@@
Specifically, in our experiments comparing with trlX, we have the following observations:
@@
Figure 9: Time cost between Separation strategy and trlX in generation and training.
@@
Flattening vs trlX Our Flattening strategy can achieve a speedup ratio up to $\times3$ compared with the trlX (with or without gradient checkpointing enabled). Gradient checkpointing[5] refers to whether activation value recalculations are enabled during training. We conduct the experiments for the 7B and 13B models on a single machine with 8 GPUs.
@@
As shown in Table IV, our Flattening strategy without gradient checkpointing accelerated the throughput from 6.04 to 10.48 samples per second, resulting in a speedup of 73.51% for training the 6B models. This improvement arises due to the single-device bottleneck in trlX, where one device is responsible for aggregating data from other devices, incurring significant communication costs and leaving other devices idle, as discussed in Section IV. Moreover, the remaining memory that is not utilized by the Reward model is wasted, limiting the overall batch size. Also, this design leads to a longer generation duration, leaving other devices idle and resulting in underutilized resources, regardless of whether trlX is in standalone or coexisted mode. It is noteworthy that trlX narrows the performance gap when the reward model can fully utilize the single device in 13B model training. Additionally, when gradient checkpointing is enabled, the performance gap increases significantly, resulting in a speedup ratio of 332%. This is because gradient checkpointing can reduce memory consumption and in turn, increase the overall training batch size. However, the single-device bottleneck in trlX disables this benefit, hindering the potential performance improvement from gradient checkpointing.
@@
Interleaving vs trlX. Our Interleaving strategy demonstrates significant improvements in training speed compared to trlX, which achieves acceleration ranging from 3.4×3.4$\times$3.4 × to $\times1$. As shown in Table V, our Interleaving strategy increases the throughput from 6.21 to 27.27 samples per second, resulting in a speedup ratio of 3.4$\times$ compared to trlX for the 7B model training on 1$\times$8 GPUs. This speedup ratio further increases to 6.1$\times$ and 11$\times$ when scaling out to 2$\times$8 and 4$\times$8 GPU devices respectively. Additionally, our Interleaving strategy improves the throughput by 6.6%, 8.3%, and 7% respectively when compared to our Flattening strategy, as we scale out to more GPUs. The training with the 13B models exhibits a similar pattern. It is analyzed that the single-GPU bottleneck in the trlX approach significantly limited the throughput, even using the maximum batch size before running out of memory under different resource configurations for the 7B and 13B models. Furthermore, the interleaving of the Reward model and Reference model in our Interleaving strategy reduces memory redundancy and decreases communication costs, resulting in considerable acceleration.
VI-B3 Time-cost Analysis
@@
We conduct a further analysis of the time duration for the training and generation stages. This analysis confirms that our placement strategy significantly improves training efficiency by substantially reducing the time-consuming generation stage.
@@
i) In the AC-NonShare scenario, we experiment with the Llama 33B model on 4$\times$8 GPU devices. Under DeepSpeed-chat’s Flattening strategy, the generation stage occupies the majority of the total duration, accounting for 93.31% in one step, while the training stage only takes up 6.69% as illustrated in Figure 9. However, our Separation strategy reduces the overall duration in one step from around 400 seconds to 30 seconds by significantly reducing the time-consuming generation stage. The percentage of the generation stage is reduced from 93.31% to 56%, almost half of the total duration. This reduction is particularly significant. It is attributed to our Separation strategy, which allows for intra-node tensor parallelism and other inference-optimized techniques for the inference models.
@@
ii) In the AC-Share scenario, we examine the time expenses for the Llama 7B model on 2$\times$8 GPU devices. As depicted in Figure.9, our separation approach can reduce the overall duration from around 72 seconds to 10 seconds by largely decreasing the generation stage. The percentage of generation duration drops from 97.86% to 34.83%. In the generation phase, the trlX approach suffers from the aforementioned single-device bottleneck. In contrast, our Separation strategy can speed up the generation stage using various optimization techniques, resulting in significant reductions in overall training time.
VII Conclusion
@@
In this paper, we introduce a novel RLHF training framework that adaptively places multiple large language models in the RLHF pipeline according to different training scenarios in terms of model and device scales. Our two placement strategies, Interleaving and Separation, not only significantly enhance training efficiency but also offer adaptable solutions for different training scenarios. Additionally, our framework is easy to use and provides automatic hyperparameter tuning for better efficiency. In the future, we plan to further improve our framework by enhancing the Planner component, such as exploring advanced techniques beyond prior or grid-based methods.

### 25 Jan 2024 v2

@@ V-B1 Comparing with DeepSpeed-Chat
Separation vs Flattening Strategy. The experiments for the Separation strategy are specifically conducted in large-scale training scenarios, such as training 33B and 65B models on $4\times 8$ and $16\times 8$ GPU devices. This is because this strategy is not well-suited for small training scenarios with a limited number of GPUs and smaller models. The reason is that the Separation strategy involves adding shadow Actor and Critic models exclusively for the prediction stage, which results in increased memory consumption.
However, the benefits of the Separation strategy are significant, as demonstrated in Table II. Our Separation strategy outperforms the Flattening strategy used in DeepSpeed-Chat by factors ranging from 4$\times$ to 11$\times$. For instance, in 33B model training using 4$\times$8 GPU devices, the throughput increases from 0.93 to 4.76 samples per second, resulting in a speedup ratio of 4$\times$. The speedup ratio further increases to 7$\times$ when the model size is increased to 65B. Additionally, the Separation strategy achieves even higher speedup when scaling out to $16\times 8$ GPU devices, reaching a speedup ratio of 11$\times$ compared to the baseline using DeepSpeed-Chat. The contribution comes from that the Separation strategy decouples the training and inference runtime of the Actor model, which is not possible for the Flattening strategy. This decoupling allows the usage of intra-node tensor parallelism for the generation stages as discussed in section IV-A. For example, in our training of 65B model using $16\times 8$ GPUs, $3\times 8$ devices are exclusively allocated for the Shadow Actor model. In this case, we apply tensor parallelism within the same node while replicating it across three machines. This approach significantly reduces inter-node communications compared to frequent inter-node AllGather and Scatter communications in ZeRO-3, which are required in the 33B and 65B model training for the Flattening strategy.
@@
Heterogeneous Separation vs DeepSpeed-Chat. We also evaluate the Heterogeneous Separation strategy in a heterogeneous GPU cluster. The cluster consists of 8 A100 GPUs and an additional 8 spare V100 GPUs with 32GB memory. These heterogeneous V100 GPUs can not be utilized under the Flattening strategy. We specifically conduct experiments using large models, such as the 33B model. Under our heterogeneous Separation strategy, the Ref and Reward models are placed on a separate node with 4 V100 GPUs respectively. We enable the intra-node tensor parallelism for both models and allocate two A100 GPUs to two Shadow Actor replicas exclusively. This configuration aims to accelerate the generation speed and improve the overall training speed.
As shown in Table III, introducing extra V100 devices under the heterogeneous Separation strategy results in a speedup of 14.49%. This acceleration increases the throughput from 0.69 to 0.79 samples per second. Although the acceleration may not be considerable due to the computing power gap between A100 and V100, this strategy explores a novel solution that allows RLHF training to utilize extra heterogeneous resources to enhance the overall throughput. It re-balances the inference workloads among different GPU devices by simply adjusting the batch size, without requiring extensive consideration of software and hardware heterogeneity among different GPUs.
V-B2 Comparing with trlX
In this subsection, we compare our APP framework against trlX, which employs a variant of the Flattening strategy, in the AC-Share scenario. We evaluate our implementation of Interleaving strategies in the APP framework in terms of throughput under various computing resources. However, before presenting the experimental results, there are three important notes to mention:
i) To align the number of models with trlX for a fair comparison, our Interleaving strategy only considers the Actor, Ref, and Reward models, excluding the Critic model in terms of placement. The reason is that the Actor and Critic models share most of the parameters in AC-Share structure. The modifications required are minimal, thanks to the modular design of our model placement abstraction.
@@ V-B2 Comparing with trlX
TABLE IV: Comparing our Interleaving strategy with trlX on $1$ machine with $8$ GPU devices. Two kinds of trlX implementation are employed.
@@
Interleaving 11.17 2.96
ON trlX (standalone RM) 5.92 2.87
@@
Interleaving 27.27 11.76
TABLE V: Comparing our Interleaving strategy with trlX on varying numbers of machines, each with $8$ GPU devices.
@@
Figure 8: Time cost between Interleaving strategy and trlX in generation and training.
@@
Figure 8: Time cost between Interleaving strategy and trlX in generation and training.
Figure 9: Comparison of maximum available batch size in AC-NonShare and AC-Share scenarios.
@@
Specifically, our Interleaving strategy can achieve a speedup ratio of up to 3.6×3.6$\times$3.6 × compared with the trlX, especially with gradient checkpointing enabled  [^4] Gradient checkpointing[4] refers to whether activation value recalculations are enabled during training. and more participating nodes.
@@
Comparison Under Gradient Checkpointing. As shown in Table IV, our Interleaving strategy without gradient checkpointing accelerated the throughput from 6.04 to 11.17 samples per second, resulting in a speedup of 84.93% for training the 6B models. This improvement arises by avoiding the single-device bottleneck in trlX, where one device is responsible for aggregating data from other devices, incurring significant communication costs and leaving other devices idle, as discussed in Section III-A. Moreover, the remaining memory that is not utilized by the Reward model is wasted, limiting the overall batch size. Also, this design leads to a longer generation duration, leaving other devices idle and resulting in underutilized resources, regardless of whether trlX is in standalone or coexisted mode. It is noteworthy that trlX narrows the performance gap when the Reward model can fully utilize the single device in 13B model training. Additionally, when gradient checkpointing is enabled, the performance gap increases significantly, resulting in a speedup ratio of 360.64%. This is because gradient checkpointing can reduce memory consumption and in turn, increase the overall training batch size. However, the single-device bottleneck in trlX disables this benefit, hindering the potential performance improvement from gradient checkpointing.
@@
Comparison When Scaling Out. Our Interleaving strategy demonstrates significant improvements in training speed compared to trlX, which achieves acceleration ranging from 3.4×3.4$\times$3.4 × to $\times1$, when scaling out. As shown in Table V, our Interleaving strategy increases the throughput from 6.21 to 27.27 samples per second, resulting in a speedup ratio of 3.4$\times$ compared to trlX for the 7B model training on 1$\times$8 GPUs. This speedup ratio further increases to 6.1$\times$ and 11$\times$ when scaling out to 2$\times$8 and 4$\times$8 GPU devices respectively. The training with the 13B models exhibits a similar pattern. It is analyzed that the single-GPU bottleneck in the trlX approach significantly limited the throughput. Furthermore, the interleaving of the Reward model and Ref model in our Interleaving strategy reduces memory redundancy as discussed in Section IV-A, resulting in considerable acceleration.
V-B3 Detailed Analysis
@@
To gain a deeper understanding of the performance enhancements, we conduct further analysis of the time cost for the training and generation stages as well as the maximum batch size available for each strategy. This analysis confirms that our placement strategy significantly improves training efficiency by substantially reducing the time-consuming generation stage and enlarging the maximum batch size.
@@
i) In the AC-NonShare scenario, we experiment with the Llama 33B model on 4$\times$8 GPU devices. Under DeepSpeed-chat’s Flattening strategy, the generation stage occupies the majority of the total duration, accounting for 93.31% in one step, while the training stage only takes up 6.69% as illustrated in Figure 9. However, our Separation strategy reduces the overall duration in one step from around 400 seconds to 30 seconds by significantly reducing the time-consuming generation stage. The percentage of the generation stage is reduced from 93.31% to 56%, almost half of the total duration. This reduction is attributed to our Separation strategy, which allows for intra-node tensor parallelism and other inference-optimized techniques for the inference models. Additionally, our Interleaving strategy increases 50% of the maximum batch size available, from 24 to 36 compared with the DeepSpeed-Chat in the 7B model, as displayed in Figure 9.
@@
ii) In the AC-Share scenario, we examine the time expenses for the Llama 7B model on 2$\times$8 GPU devices. As depicted in Figure 9, our approach can reduce the overall duration from around 72 seconds to 10 seconds by largely decreasing the generation stage, as the percentage of generation duration drops from 97.86% to 34.83%. In the generation phase, the trlX approach suffers from the aforementioned single-device bottleneck. In contrast, our Interleaving strategy can speed up the generation stage by placing the Reward model on multiple devices and interleaving it with the Ref model. For the maximum batch size available, the Interleaving strategy increases the batch size 25$\times$ compared with trlX in the 13B model, rocketing up from 2 to 50, as illustrated in Figure 9.
VI Conclusion
@@
In this paper, we introduce a novel RLHF training framework that adaptively places multiple large language models in the RLHF pipeline according to different training scenarios in terms of model and device scales. Our two placement strategies, Interleaving and Separation, not only significantly enhance training efficiency but also offer adaptable solutions for different training scenarios. Additionally, our framework is easy to use and provides simple user interface.